# Generative AI for Physical-Layer Authentication

### v2

@@ VI Conclusions
In this paper, we have first revisited classic identity authentication techniques, including cryptography-based authentication, blockchain-based authentication, PLKA, and PLA. Then, we have reviewed representative GAI models, including GANs, VAEs, GDMs, and Transformers, and we have analyzed the limitations of applying DAI in PLA and have pointed out the potential of GAI-based PLA. Furthermore, we have provided GAI-enhance methods for PLA in detail, including fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission in fingerprint collection phase, attack detection, multiuser authentication, and efficient training and deployment in model training phase, and flexibility for eliminated legitimate users, defense against intelligent attacks, and adaptation to different scenarios in performance optimization phase. Additionally, we have presented a case study to show the superiority of GAI over DAI in authentication performance under dynamic environments. Lastly, we have provided challenges and future directions.

### v3

@@ VI Conclusions
In this paper, we first provide a concise overview of identity authentication techniques and GAI models. We then analyze the limitations of applying DAI in PLA and explore how GAI can address these challenges. Furthermore, we propose a structured taxonomy for GAI-enhanced PLA methods, delineating three key stages of the PLA pipeline: fingerprint collection, model training, and performance optimization. Specifically, GAI contributes to (1) fingerprint collection: enhancing accuracy, enriching diversity, ensuring completeness, and improving transmission efficiency of collected fingerprints; (2) model training: tailoring models to meet specific task requirements; and (3) performance optimization: increasing flexibility to accommodate legitimate users, strengthening robustness against intelligent adversarial attacks, and boosting adaptability across diverse operational scenarios. Additionally, we introduce a novel PLA framework leveraging channel extrapolation and GAI, supported by a case study using GDMs to demonstrate GAI’s superiority in enhancing PLA robustness. Finally, we outline a forward-looking research roadmap for GAI-based PLA.